OpenStudy (anonymous):

interval and set notation for X<3

OpenStudy (anonymous):

Interval : \[] - \infty ; 3[ \] Set builder : \[\left \{ x | x \in \mathbb{R}, x \in ] - \infty ; 3[\right \}\]

OpenStudy (anonymous):

Interval:\[(-\infty,3)\] Set Builder:\[\{x| (x \in \mathbb{R}) \wedge (x < 3)\}\]